# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v2

@@ 7 Conclusion and Outlook
This paper has provided a systematic overview of text masking approaches for VLM training, which accelerate training and have the potential to also improve performance. In this paper, we demonstrate that training epochs have a substantial impact on the performance of different text masking strategies. We conclude, contrary to the existing state-of-the-art, CLIPA [12], that syntax masking is not the best performing text masking approach and that truncation masking can be avoided. We carry out an analysis of word frequency distributions that shows that frequency is what you need to think about when text masking the training data of VLMs. Specifically, both syntax and truncation have visible effects on the word frequency distribution, making clear why they do not perform well.
@@
We have introduced CLIPF, an approach for text masking based on word frequency that improves the performance of CLIP. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which is time-consuming to process. CLIPF is also interesting because it controls frequency directly and in doing so seems to balance nouns with words in the “Other” POS categories, in contrast to other approaches.
@@
We carry out an extensive comparison among existing text matching strategies (CLIPA variants) and CLIPF that revals the strong performance of CLIPF in a variety of zero-shot classification tasks and when scaling to larger data sets. CLIPF demonstrates a notable ability to rapidly learn image and text representations when compared to a model that pre-trains on the full text and other text masking strategies. It also delivers strong performance in under highly aggressive masking (i.e., using only 4 tokens per training text). Moving forward, the insights gained from this paper about the importance of frequency can be used to further optimize CLIPF to directly optimize the word distribution allowing for more tuning control when trading off VLM performance and VLM training speech during text masking.
Appendix A Text Masking Analysis
A.1 Text Masking Cases
Figure 5: The curve of Formula 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of formular 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$,$10^{-8}$.
Table  9 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in Fig. 3, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content.
Table 10 shows the masking probabilities for certain words. High-frequency words such as “of" “the" “and" and “." are highly likely to be masked from the text.
@@ A.1 Text Masking Cases
Table 10: The probability of masking words in the text is calculated using the formula provided in Equation 2. The example is selected from CC12M [15]. The value of $t$ of formular 2 is set to $10^{-6}$.
@@
A.2 Word Category Distribution Across Different Text Masking Strategies
As shown in Table 11, we provide the word type counts of Table 5. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
Masking NN Count JJ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
@@ A.2 Word Category Distribution Across Different Text Masking Strategies
Figure  6 shows the word cloud before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying frequency word masking, the word cloud in Table 6(b) shows a more balanced distribution of word frequencies compared to before the masking. The word “the" appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full-text. As depicted in Table 6(f), syntax masking tends to retain more nouns such as “person," “stock," and “illustration" that show a different pattern from the others.
Fig. 6 shows word clouds before and after applying text masking. In these visuals, the font size represents the relative frequency of each word. Words such as “the,” “of,” and “person” appear larger than others, indicating their higher frequency. After applying word frequency masking, the word cloud in Fig. 6(b) displays a more balanced distribution of word frequencies compared to before masking. The word “the” appears smaller than in other word clouds. The reduction in frequent words allows infrequent words to become more prominent. The truncation, random, and block masking techniques result in word cloud patterns similar to those seen with the full text. As depicted in Fig. 6(f), syntax masking tends to retain more nouns such as “person”, “stock”, and “illustration”, resulting in a different pattern from the others.
Refer to caption Refer to caption
(a) Before masking (b) Frequency
Refer to caption Refer to caption
(c) Truncation (d) Random
Refer to caption Refer to caption
(e) Block (f) Syntax
Figure 6: The word cloud on the left displays the full text, while the one on the right shows the text after applying text masking, where we retain 6 words. The font size of the word represents the relative word frequency for each figure. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.3 Word Distribution Analysis for CLIPF
@@
Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of formula 2 is set to $10^{-6}$.
A.4 Word Distribution of CLIPF and SW-CLIP
We also compare the word distributions of CLIPF and SW-CLIP after applying text masking. As illustrated in Fig. 8, SW-CLIP masks more frequent words based on its threshold, preventing it from fully utilizing the input slots of the text encoder. Consequently, SW-CLIP exhibits a flatter word distribution compared to CLIPF.
Figure 8: The figure illustrates the distribution of top-50 words in the text for SW-CLIP and CLIPF.
Appendix B More Results
B.1 The details of Image-text Retrieval
In Table 12, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
Models Masking Text Tokens Text Retrieval Image Retrieval
@@ B.1 The details of Image-text Retrieval
CLIP 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
FLIP 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
CLIPA truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
random 16 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
block 16 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
syntax 16 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
CLIPF frequency 16 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
CLIPA truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
random 8 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
block 8 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
syntax 8 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
CLIPF frequency 8 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
CLIPA truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
random 6 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
block 6 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
syntax 6 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
CLIPF frequency 6 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
CLIPA truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
random 4 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
block 4 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
syntax 4 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
CLIPF frequency 4 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
B.2 Learn Curved for Image-text Retrieval
We also investigate the zero-shot image-text retrieval performance over training epochs on the COCO and Flickr30k datasets. It also shows that the number of training epochs has a substantial impact on the performance, as illustrated in Fig. 9. And the performance gap between syntax masking and the other text masking on the zero-shot image-text retrieval task is larger than zero-shot classification task. This is because the zero-shot image-text retrieval task requires the model not only to understand the objects/nouns, but also to understand the logic of the text. Moreover, we present the zero-shot image-text retrieval performance across training epochs for models pre-trained on the LAION-400M dataset. As illustrated in Fig. 10, the performance patterns of models trained on LAION-400M are similar to those of previous models that pre-trained on the CC12M dataset. Specifically, CLIPF consistently outperforms syntax masking. Block masking achieves comparable performance to syntax masking and even surpasses it on the image-to-text retrieval task using the COCO dataset at epoch 16th.
Figure 9: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model for an additional epoch without image and text masking.
Figure 10: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on LAION-400M for 16 epochs with 25 image tokens and 4 text tokens to speed up training and fine-tune the model for 0.4 epoch with 224x224 image size and 32 text tokens.
Additionally, Fig. 11 illustrates the validation loss across pre-training epochs. It is evident that the validation loss for truncation and syntax masking methods begins to increase earlier compared to other text masking methods, suggesting that these methods may overfit to the data sooner.
Figure 11: The validation loss curve of COCO and Flick30k datasets over the pre-training epochs.
B.3 Learn Curved for DTD and OxfordPets Datasets
The learning curves for the DTD [40] and OxfordPets [41] datasets are presented because not all class names in these datasets are nouns. As shown in Table 13, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in Fig. 12. Moreover, the performance of syntax masking is very poor and almost close to truncation masking. OxfordPets is a dataset containing incomplete noun class names. In the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking. However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
Figure 12: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
B.4 Learn Curved before fine-tuning
In Fig. 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning in Fig. 15 (30 epochs) and Fig. 16 (45 epochs). They have a similar pattern to the models before fine-tuning for every text masking strategy.
Figure 15: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training.
Figure 16: Zero-shot classification accuracy on ImageNet-1K over the 45 epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 45 epochs with image and text masking (75%) to speed up training.
B.5 Fine-tuning the models with More Epochs
The purpose of fine-tuning the models without image and text masking is to reduce the distribution gap between pre-training and evaluation. As demonstrated in FLIP and CLIPA, a small number of fine-tuning steps is usually sufficient to bridge this gap [8, 12]. We conducted additional experiments using two fine-tuning epochs, shown in Tab 13, to demonstrate that increasing the number of fine-tuning epochs does not substantially improve performance but instead leads to higher computational costs.
Table 13: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC12M for 30 epochs with 75% image masking for speedup (pre-train). Then, we fine-tune the model with one epoch (fine-tune-1) and two epochs (fine-tune-2) without image and text masking.
Models Masking Image Tokens Text Tokens CC12M
pre-train fine-tune-1 fine-tune-2
CLIP 197 32 36.6
FLIP 49 32 32.0 33.7 34.1
CLIPA truncation 49 8 25.4 28.4 28.5
random 34.5 36.9 37.1
block 35.5 37.9 38.0
syntax 28.5 35.0 35.5
CLIPF frequency 36.6 39.3 39.4
B.6 Comparison with SW-CLIP
In this section, we provide a comparison of CLIPF with SW-CLIP [22], which also uses frequency-based sampling but imposes a threshold on the frequency score. The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold. In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots. For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [22]. It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer. For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking. Subsequently, we fine-tuned both models with an additional epoch without any image or text masking. The results in Table 14 show that after fine-tuning CLIPF outperforms SW-CLIP.
@@ B.6 Comparison with SW-CLIP
B.7 Applying text masking on SigLIP

### v3

@@ 7 Conclusion and Outlook
This paper has provided a systematic overview of text masking approaches for VLM training, which can boost performance and reduce computation. VLMs are important because of their wide use in practical applications, including improving Multimodal Large Language Models (MLLMs) [18, 25]. We have carried out an analysis of how text masking impacts word frequency in the training data and discovered that text masking strategies perform better if they do not disproportionately impact word frequency. Our experiments also demonstrate that training epochs have a substantial impact on the performance of different text masking strategies. We conclude, contrary to the existing state-of-the-art, CLIPA [20], that syntax masking is not the best performing text masking approach.
@@
Further, this paper has introduced CLIPF, a strategy for text masking based on word frequency. Our experiment shows that CLIPF yields strong performance and shows the ability to outperform the original CLIP, while using less training resources. Unlike syntax-based masking methods, CLIPF does not require part-of-speech (POS) information, which consumes more resources to generate than frequency statistics. CLIPF is also interesting because it seems to balance nouns with words in the “Other” POS category, in contrast to syntax masking.
@@
We carry out a quantitative comparison among CLIPF and existing text matching strategies that reveals the strong performance of CLIPF in a large range of zero-shot classification tasks. CLIPF demonstrates a notable ability to learn image and text representations when compared to a model pre-trained on text masked with other existing text masking strategies. It also outperforms other approaches under highly aggressive masking (i.e., using only 4 tokens per training text).
This paper has argued that frequency is what you need to think about when choosing a text masking strategy. Specifically, both truncation and syntax masking have a disproportionate impact on the frequencies of specific words, making clear why they do not perform well. Further, CLIPF, which exploits word-frequency directly, is a strong performer. These insights are a valuable guide for researchers and developers, who need to select an appropriate text masking strategy, while taking in to account available computational resources. Moving forward, CLIPF can be further optimized to more directly shape the word frequency distribution, providing even more control when trading off VLM performance and computation during text masking.
Appendix A Details of Experimental Setup
Dataset Data set statistics are summarized in  Tab. 7. For the majority of our experiments, we pre-train the models on Conceptual Captions 3M (CC3M) [36], Conceptual 12M (CC12M) [3]. These datasets have been used by OpenCLIP, DeCLIP, SLIP, A-CLIP and BLIP [15, 21, 30, 43, 18] To reproduce CLIPA, we use LAION-400M [35], which is used by CLIPA. LAION-400M is approximately 133 times larger than CC3M and 32 times larger than CC12M, indicating that it contains a broader range of concepts and offers greater diversity compared to CC3M and CC12M.
Note that the difference in caption length with have an impact on masking performance, although we do not directly measure the difference in this paper.
Architecture For the image encoder, we use ViT-B/16 (86M parameters) [9] architectures with global average pooling and learnable positional embeddings. For the text encoder, we use a Transformer-based model [40] and byte-pair encoding with a 49K token vocabulary. Additionally, we run one experiment with a larger image encoder (VIT-L/16) with 303M parameters and report the results in the Appendix C of supplementary material. We chose these encoders because they are the largest ones available to us, given our current resource constraints. The input image size is 224 for all datasets. When using full text for training, the maximum context length is 32. Zero-padding is applied to input text that is shorter than the maximum token length of the model.
Dataset Samples Total words Caption length
CC3M $2.72\times 10^{6}$ $2.80\times 10^{7}$ $10.30\pm 4.72$
CC12M $9.30\times 10^{6}$ $06\times 10^{8} 22.$$2.15\pm 17.20$
LIAON-400M $2.98\times 10^{8}$ $3.71\times 10^{9}$ $12.51\pm 15.82$
Config Pre-training Fine-tuning
optimizer AdamW [27] AdamW [27]
learning rate 1e-3 1e-5
weight decay 0.2 0.2
learning rate schedule cosine decay [26] cosine decay [26]
warmup steps 10k 10%
epoch 30 1
$t$ $10^{-6}$
$\tau$ 0.07
numerical precision amp amp
RandomResizedCrop (50, 100) (50, 100)
Table 8: Details of the pre-training and fine-tuning setups.
Image masking Text masking Image tokens Text tokens Total Percentage
0% 0.00% 196 32 228 100%
75% 0.00% 49 32 81 35.53%
50.00% 49 16 65 28.51%
75.00% 49 8 57 25.00%
81.25% 49 6 55 24.12%
87.50% 49 4 53 23.25%
Training Following CLIP, OpenCLIP [32, 15], we pre-train our model for 30 epochs on the CC3M and CC12M datasets. For the LAION-400M dataset, we pre-train the model for 16 epochs, extending CLIPA’s experiments, which were conducted for 6 epochs on the same dataset. Details of the pre-training configuration are given in Tab. 8.
[…]
During text masking, we reduce the number of tokens from 32 to 16, 8, 6, and 4. To measure the training speed of CLIP, CLIPA, and CLIPF, we compare the number of text tokens processed by each model. As shown in  Tab. 9, the number of text tokens processed by each model during pre-training is calculated based on different image and text masking ratios. When we pre-trained the model using image masking, the total number of tokens is 81 and the percentage of text tokens is 39.5%. When we apply 50% text masking, the total number of tokens is 65. Compared to training without text masking, this results in a speed increase of approximately 20%. Moreover, when we apply 87.5% text masking, the total number of tokens is 53, resulting in a speed increase of approximately 34% compared to training without text masking.
Fine-tuning Following FLIP, and CLIPA, in order to bridge the distribution gap between pre-training and evaluation, we fine-tuned the model without images and text masking. Note that, fine-tuning the models without masking in masking work is focused on reducing the distribution gap and applied sparingly to avoid undoing the tradeoff. The details of the fine-tuning configuration are provided in  Tab. 8.
Figure 5: The curve of Equation 2. The x-axis is the word frequency $f(w_{i})$, and the y-axis is the $P(w_{i})$. The value of $t$ of Equation 2 is set to $10^{-5}$, $10^{-6}$, $10^{-7}$$10^{-8}$.
@@ Appendix A Details of Experimental Setup
Table 11: The probability of masking words in the text is calculated using the formula provided in Equation 2. The example is selected from CC12M [3]. The value of $t$ of Equation 2 is set to $10^{-6}$.
Figure 6: The complete x-axis label for Fig. 3. in Sec. 3 We set the text length after text masking to 6. The x-axis represents the word index, which is sorted by counts of the original data, and the y-axis shows the word frequency. The dataset used is CC12M and the value of $t$ of Equation 2 is set to $10^{-6}$. We remove special characters from the vocabulary.
Masking NN Count $J$ VB OTHER NN (%) JJ (%) VB (%) OTHER (%) Total
@@
Appendix B Text Masking Analysis
B.1 Text Masking Cases
Tab. 10 presents the potential text resulting from the text masking technique. Since truncation is fixed in each epoch, it may result in the loss of important information at the end of the text. Certain words tend to appear in specific positions within the text; for example, “the” and “a” are most likely to be the first words of a sentence. As shown in  Fig. 6, truncation retains more occurrences of “the” and “a” than other text masking strategies. In contrast, random and block strategies can generate different texts in each epoch for text data augmentation, but they may retain some words with a high frequency. Syntax masking retains the nouns in the text and remains the same in each epoch. However, this strategy may cause the model to overfit on the frequency of noun words. In contrast, CLIPF varies the text in each epoch, as words may be retained or removed according to their frequency. This strategy serves two primary purposes: it enhances text diversity and reduces the risk of overfitting to frequent words. Another advantage of CLIPF is that it can remove frequent prepositions that are less directly relevant to the objects in the image, such as “a,” “in,” “of,” and others. This helps the model focus on the most helpful aspects of the content.  Tab. 11 shows the masking probabilities for certain words. High-frequency words such as “of” “the” “and” and “.” are highly likely to be masked from the text.
B.2 Word Category Distribution Across Different Text Masking Strategies
As shown in  Tab. 12, we present the word type counts corresponding to Tab. 1 in the main paper. After applying text masking strategies such as truncation, random, block, syntax, and CLIPF, the word counts remain similar. However, SW-CLIP does not maximize the use of the input slots, utilizing only 85% of the words compared to other text masking strategies. Additionally, SW-CLIP masks a high percentage of other types of words, which may impact the model’s zero-shot ability.
Figure 7: Frequency-based text masking strategies vary according to the number of text words used during pre-training. The dataset used is CC12M and the value of $t$ of Equation 2 is set to $10^{-6}$.
Models Masking Image Tokens Text Tokens ViT-B/16
IN-A IN-O IN-R IN-S IN-V2 ON
CLIP 197 32 8.97 37.85 49.11 25.70 31.48 24.20
CLIPF frequency 98 8 10.37 37.75 52.33 28.52 35.62 24.00
Table 13: Zero-shot robustness evaluation. Comparison of the zero-shot accuracy performance of CLIP and CLIPF on various datasets when using more image tokens. The models are pre-trained on CC12M [3] for 30 epochs with image masking (50%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
Flickr30k COCO Flickr30k COCO
R@1 R@1 R@1 R@1
CLIP 197 32 62.62 35.54 45.42 24.22
CLIPF frequency 98 8 63.11 37.14 46.84 24.89
Table 14: Zero-shot Image-Text retrieval evaluation. We evaluated CLIP, CLIPF image-text retrieval performance on the COCO and Flickr30k datasets when using more image tokens. The models are pre-trained on CC12M [3] for 30 epochs with image masking (50%) to speed up training and fine-tune the model an additional epoch without image and text masking.
B.3 Word Distribution Analysis for CLIPF
@@
Models Masking Image Tokens Text Tokens Text Retrieval Image Retrieval
@@
CLIP 197 32 62.62 86.00 91.81 35.54 62.38 74.08 45.42 72.56 81.50 24.22 48.42 60.42
FLIP 49 32 54.73 80.37 87.97 29.34 56.08 67.00 38.62 66.09 75.40 20.88 43.22 54.78
FLIP truncation 16 44.67 73.08 81.85 25.54 51.90 64.64 34.99 61.05 70.85 19.64 41.91 53.51
random 58.48 84.62 90.53 32.36 58.76 69.98 43.61 70.67 80.04 22.81 46.00 57.71
CLIPA block 56.51 81.26 89.05 30.82 58.32 70.38 44.06 71.20 80.10 22.66 46.24 58.06
syntax 54.54 81.07 88.36 29.60 56.52 68.54 41.16 68.32 77.51 21.40 44.82 56.56
CLIPF frequency 57.89 84.62 90.04 31.52 58.38 70.30 42.72 69.61 78.60 22.57 46.15 57.95
FLIP truncation 8 30.47 59.47 70.51 16.92 38.62 51.34 23.96 47.29 57.85 12.78 31.66 42.92
random 58.58 84.52 91.81 36.24 62.16 72.90 43.79 70.89 80.14 23.16 46.74 58.86
CLIPA block 60.06 85.01 92.11 35.88 62.58 73.74 45.98 73.23 82.49 24.65 48.81 60.70
syntax 50.30 78.30 86.98 29.64 55.42 67.18 38.46 66.77 77.36 20.28 43.78 55.40
CLIPF frequency 58.68 85.10 91.51 34.74 61.38 71.88 44.57 72.58 81.92 23.30 47.50 59.24
FLIP truncation 6 32.25 61.05 72.98 16.18 39.52 52.48 22.39 47.46 59.39 11.93 30.03 40.86
random 55.23 82.54 89.25 32.58 57.54 68.68 42.47 70.12 79.43 21.34 44.79 56.26
CLIPA block 54.44 79.68 88.66 33.12 58.96 70.64 41.68 69.13 79.33 21.73 45.23 57.29
syntax 46.15 76.04 84.81 26.78 52.56 64.62 34.08 61.76 73.25 17.99 40.22 52.05
CLIPF frequency 56.02 82.05 88.56 32.32 58.58 70.00 41.05 69.17 78.88 21.28 44.55 56.05
FLIP truncation 4 30.97 56.51 67.06 17.44 38.48 49.76 21.09 43.77 55.70 11.07 27.98 38.54
random 41.62 69.53 80.18 24.14 48.32 60.14 30.26 56.92 68.42 15.63 35.72 47.30
CLIPA block 40.83 69.03 79.09 23.56 47.94 59.20 30.77 57.85 69.29 14.98 35.64 47.35
syntax 38.66 65.88 75.74 21.68 44.38 56.60 26.08 52.80 64.32 14.14 33.71 44.98
CLIPF frequency 45.07 72.49 81.95 25.60 49.98 61.96 31.72 59.35 71.66 15.96 36.58 48.32
@@
Appendix C More Results
C.1 More image tokens
As shown in Tab. 13 and Tab. 14, CLIPF achieves better zero-shot robustness in 4 out of 6 datasets and image–text retrieval performance in both Flick30k and COCO datasets than the original CLIP, even though it only uses 50% of the image tokens and 25% of the text tokens.
C.2 The details of Image-text Retrieval
In Tab. 15, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
Figure 9: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies after fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens ViT-L/16
pre-train fine-tune
FLIP truncation 49 8 28.4 31.7
random 36.0 38.3
CLIPA block 36.8 39.5
syntax 32.4 37.5
CLIPF frequency 38.0 40.2
Table 17: Comparison of the ImageNet-1k zero-shot accuracy performance of ViT-L/16 with different text masking strategies. The models are pre-trained on CC12M [3] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
C.3 Learn Curved for DTD and OxfordPets Datasets
The learning curves for the DTD [6] and OxfordPets [31] datasets are presented because not all class names in these datasets are nouns. As shown in Fig. 8, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in  Fig. 9. Moreover, the performance of syntax masking is very poor and almost close to truncation masking. OxfordPets is a dataset containing incomplete noun class names. As shown in Fig. 10, in the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking. However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
C.4 Comparison with SW-CLIP
In this section, we provide a comparison of CLIPF with SW-CLIP [23], which also uses frequency-based sampling but imposes a threshold on the frequency score. The effect of this threshold is that input tokens will go unused if not enough words in a training text surpass the threshold. In contrast, CLIPF calculates the word frequency by the threshold to prioritize words and maximize the input slots. For instance, using CC3M as an example, the average length of text before masking is 10.31 ± 4.7, and after SW-CLIPF masking, it is 4.26 ± 2.6 [23]. It is clear that the use of the input slots is not maximized when the input length of text tokens is set to 6 or longer. For a fair comparison with CLIPF, we pre-train SW-CLIP using the same setup: 75% image masking and 81.25% text masking. Subsequently, we fine-tuned both models with an additional epoch without any image or text masking. The results in Tab. 16 show that after fine-tuning CLIPF outperforms SW-CLIP.
C.5 Large Image Encoder Architecture
We carried out an extra experiment which is apply different text masking strategies on larger architectures which is ViT-L/16. The behavior of different architectures is very consistent when applying different text masking strategies, as shown in Tab. 17.
C.6 Applying text masking on SigLIP